# PromptFix: You Prompt and We Fix the Photo

### v1

@@ Abstract
Diffusion models equipped with language models demonstrate excellent controllability in image generation tasks, allowing image processing to adhere to human instructions. However, the lack of diverse instruction-following data hampers the development of models that effectively recognize and execute user-customized instructions, particularly in low-level tasks. Moreover, the stochastic nature of the diffusion process leads to deficiencies in image generation or editing tasks that require the detailed preservation of the generated images. To address these limitations, we propose PromptFix, a comprehensive framework that enables diffusion models to follow human instructions to perform a wide variety of image-processing tasks. First, we construct a large-scale instruction-following dataset that covers comprehensive image-processing tasks, including low-level tasks, image editing, and object creation. Next, we propose a high-frequency guidance sampling method to explicitly control the denoising process and preserve high-frequency details in unprocessed areas. Finally, we design an auxiliary prompting adapter, utilizing Vision-Language Models (VLMs) to enhance text prompts and improve the model’s task generalization. Experimental results show that PromptFix outperforms previous methods in various image-processing tasks. Our proposed model also achieves comparable inference efficiency with these baseline models and exhibits superior zero-shot capabilities in blind restoration and combination tasks. The dataset and code will be aviliable at https://github.com/yeates/PromptFix.
Figure 1: We propose PromptFix, a unified diffusion model capable of performing multiple image-processing tasks. It can understand user-customized editing instructions and perform the corresponding tasks. One of the key advantages of PromptFix is high-frequency information reservation, as demonstrated in tasks such as low-light enhancement, ensuring which image details are maintained throughout the VAE reconstruction process.
@@ 1 Introduction
In recent years, diffusion models [55, 17, 61] have achieved remarkable advancements in text-to-image generation. Benefiting from large-scale training on image-text pairs [56], these models can generate highly realistic and diverse images that align with text prompts. They have been successfully applied to various real-world applications, including visual design, photography, digital art, and the film industry. In addition, models trained with instruction-following data [7] have shown promising results in understanding human instruction and performing the corresponding image-processing tasks. Previous studies [21, 20, 71, 69, 19] have illustrated that with the instruction-following data, we can simply fine-tune a text-to-image generation model to perform various vision tasks such as image editing [19, 7, 71], object detection [20], segmentation [21], inpainting [69, 21], and depth estimation [20]. To follow the success of these methods, we train our model utilizing input-goal-instruction triplet data for low-level image-processing tasks.
@@
We first overcome the challenge of lacking the instruction-following data for low-level tasks. Specifically, we collect image pairs by generating degraded images from the source images and adopt data from existing datasets. Then we employ GPT4 [49] to generate the diverse text instructions for each task. We obtain $\sim$ 1.34 million input-goal-instruction triplets in the collected dataset. This dataset covers various low-level tasks including image inpainting, object creation, image dehazing, colorization, super-resolution, low light enhancement, snow removal, and watermark removal. We enrich the dataset through back-translation augmentation by swapping the inpainted and original images within the triplet and inverting the semantic orientation of the prompt. This technique effortlessly converts datasets from object removal to object creation. We also provide comprehensive details in Section 3.
@@
With the dataset, we design a new diffusion-based model named PromptFix that can understand user-customized instructions and perform the corresponding low-level image-processing tasks. In PromptFix, we address several challenges that compromise the performance of the model. First, the use of stable-diffusion architecture [55] as a generative prior often faces the issue of spatial information loss, this is caused by VAE compression [18]. Unlike unconditional or text-to-image generation, maintaining spatial detail consistency in image processing poses significant challenges, particularly with high-frequency components like text, as shown in Figure 5. To tackle this problem, we introduce the High-frequency Guidance Sampling, in which we use a low-pass filter operator [52, 48] to calculate the fidelity constraint and integrate VAE skip-connect features during inference with a lightweight LoRA [25] fusion. Second, since the generative prior is not trained on low-level images, so relying solely on instructions may not always yield the desired outcome, especially when the image degradation is severe. To tackle this degradation adaptation problem, we introduce an auxiliary prompt module to provide models with more descriptive text prompts to enhance controllability for image generation. The auxiliary text prompt can be obtained by VLMs  [43]. This approach introduces a semantic caption for a degraded image and the description of its defects, such as blurriness or insufficient lighting. The auxiliary prompt module is implemented by an additional attention layer in diffusion U-Net that adapts both instruction and auxiliary prompts as conditions and intermittently omits instructional prompts during training. We identify three key advantages of this approach: 1) enabling the model to process images with severe degradation, such as extremely low-resolution images, 2) adapting the model for blind restoration for different types of image degradation, and 3) providing additional pathways for a more precise semantic representation of the target image.
@@
  • We propose a comprehensive dataset tailored for low-level image processing tasks. The dataset contains $\sim$ 1.34 million diverse paired input-output images along with corresponding image editing instructions.
@@ 2.1 Instruction-guided Image Editing
Instruction-guided image editing significantly improves the ease and precision of visual manipulations by adhering to human directions. In traditional image editing, models primarily focused on singular tasks such as style transfer or domain adaptation [22, 54], leveraging various techniques to encode images into a manipulatable latent space, such as those used by StyleGAN [33]. Concurrently, the advent of text-to-image diffusion models [55, 24, 58, 60] has broadened the scope of image editing [7, 23]. Kim et al. [34] showed how to perform global changes, whereas Avrahami et al. [4] successfully performed local manipulations using user-provided masks for guidance. While most works that require only text (i.e., no masks) are limited to global editing [15, 36]. Bar-Tal et al. [6] proposed a text-based localized editing technique without using any mask, showing impressive results. For local image editing, precise manipulations are possible by inpainting designated areas using either user-provided or algorithmically predicted masks [14, 63, 70], all while preserving the visual integrity of the adjacent areas. In contrast, instruction-based image editing operates through direct commands like “add fireworks to the sky,” avoiding the need for detailed descriptions or regional masks. Recent approaches utilize synthetic input-goal-instruction triples [7] and incorporate human feedback [73] to execute editing instructions effectively. Despite the advances in using diffusion models for various instruction-guided image editing tasks, there is still a notable gap in research specifically addressing instruction-guided image restoration with these models. Our study aims to bridge this gap by collecting a comprehensive dataset of paired low-level instruction-driven image editing examples and proposing an all-in-one model for low-level tasks and editing.
@@ 2.2 Large Language Models for Vision
Recent advancements in the development of Large Language Models (LLMs) have led to the emergence of powerful models with extensive capabilities [43, 29, 13, 39, 74]. These LLMs, pretrained on large-scale internet-based datasets, are equipped with broad knowledge bases that enhance their zero-shot and in-context learning abilities [49, 5, 27]. Furthermore, there is a growing focus on using LLMs for multimodal tasks [3, 45, 42, 28], incorporating methods like vision-language alignment and adapter fine-tuning. These techniques ensure that the visual data processed by visual encoders is semantically aligned with the textual input of LLMs [43]. This approach has spurred significant advancements in Text-to-Image generation, leading to the development of various LLM-based diffusion models for these tasks [41, 26, 40]. Despite these successes, there remains a relative scarcity of research focusing on using large Vision-Language Models (VLMs) for instructed image editing, particularly in detailed, low-level editing tasks.
@@ 3 Data Curation
Current off-the-shelf image datasets [7, 71, 73] with instructional annotations primarily facilitate image editing research, encompassing tasks such as color transfer, object replacement, object removal, background alteration, and style transfer. Nevertheless, their overlap with low-level applications is limited. Moreover, we find it challenging to achieve satisfactory results for existing models in image restoration. Our goal is to construct a comprehensive visual instruction-following dataset specifically for low-level tasks. We obtain $\sim$ 1.34 million training triplet instances.
Paired Image Collection. We initially gather source images from various existing datasets. Subsequently, we produce degraded and inpainted images to create an extensive set of paired image data. We compile approximately two million raw data points across eight tasks: image inpainting, object creation, image dehazing, image colorization, super-resolution, low-light enhancement, snow removal, and watermark removal. For the test set, we randomly select 300 image pairs for each task. More details about the dataset composition are provided in the Appendix.
@@ 4.1 Diffusion Model
where $\alpha_{t}$ and $\sigma_{t}$ are coefficients that manage the signal-to-noise ratio at each interpolation point $\mathbf{z}_{t}$. This process aims to maintain variance, adopting coefficient strategies as detailed in sources such as [32]. Modeled as a stochastic differential equation (SDE) in continuous time, the forward process can be expressed as $d\mathbf{z}=\mathbf{f}(\mathbf{z},t)dt+g(t)d\mathbf{w_{t}},$ where $\mathbf{f}(\mathbf{z},t)$ is a vector-valued drift coefficient, $g(t)$ the diffusion coefficient, and $\mathbf{w_{t}}$ represents Brownian motion at time $t$.
@@
During the training phase, a simple diffusion loss [24] is utilized, whereby the neural network still employs forward inference to predict noise. The sample data estimate $\mathbf{z}_{0}$ can be obtained at any step $t$ by using the current noisy data and the predicted noise, and is derived as:
@@ 4.3 High-frequency Guidance Sampling
There is a fundamental requirement in image restoration and generation tasks: the processed image must maintain high accuracy in semantic. We observe that vanilla VAE reconstructions tend to lose image details such as textual rendering, which contain high-frequency information, as shown in Figure 5. Therefore we propose high-frequency guidance sampling to balance the quality and fidelity of generation.
@@
Table 1: Quantitative comparison is conducted across seven low-level datasets with a $512\times 512$ input resolution. Expert models refer to approaches, such as Diff-plugin [46], which use non-generalizable training pipelines and maintain separate pre-trained weights for each of the four restoration tasks. Image Restoration Generalist Methods denote models that integrate multiple low-level tasks into a single framework. Instruct-driven Diffusion Methods represent diffusion generative baselines that follow human language instructions. $\uparrow$ indicates higher is better, and $\downarrow$ indicates lower is better. The best and second best results are in bold and underlined respectively.
@@
We propose a fidelity constraint to model the spatial discrepancy between the image and the ground truth. We implement two types of low-pass operators to extract high-frequency signals from the degraded image. For the Fourier filtering operator $\mathcal{F}(\cdot)$, we convert the generated image from the spatial to the frequency domain using the Discrete Fourier Transform [48]. High-frequency components are then isolated via high-pass filtering and reconstituted into an image through the Inverse Fourier Transform [48]. Meanwhile, we apply the Sobel [52] edge detection operator $\mathcal{S}(\cdot)$ as a complement. The fidelity constraint evaluates the divergence between the high-frequency components of the ground truth and the processed image, ensuring the preservation of spatial information throughout the sampling process. Additionally, to obtain the image at time step $t$, we utilize predicted noise $\epsilon$ from diffusion model $\mathcal{H}$ to compute the $\mathbf{z}_{t\rightarrow 0}$ estimation at any time step. The fidelity constraint is calculated as follows:
@@
Given that $\mathbf{z}_{t>0}$ represents a noisy latent variable, assigning equal weight to each timestep’s latent is impractical. To mitigate the cumulative error induced by this practice, we introduce a time-scale weight $e^{-\lambda t}$ applied to the fidelity constraint. This exponential decay significantly reduces the impact of high noise at larger time steps. We provide a detailed discussion on the selection of the hyperparameter $\lambda$ in the Appendix. The overall sampling algorithm is detailed in Algorithm 4.1.
@@ 5.1 Experimental Setup
Baselines and metrics. We adopt instruction-based generalist models, such as InstructP2P [7], MGIE [19], and InstructDiffusion [21], as our primary comparison. MGIE employs VLM-guided techniques for image editing, while InstructDiffusion addresses overlapping tasks with our training objectives, including watermark removal and inpainting. Additionally, we evaluate all-in-one image restoration methods like AirNet [38] and PromptIR [51] (which do not support instruction input), as well as image restoration expert models fine-tuned for specific sub-tasks [46, 69]. We assess the similarity of the generated images to the ground truth using metrics such as PSNR, SSIM [65], and LPIPS [72]. For no-reference image quality evaluation, we utilize the ManIQA [68] metric.
Figure 4: Qualitative analysis of VLM-guided blind restoration for desnowing, dehazing, and low-light enhancement. The results are obtained from PromptFix without explicit task instructions, relying solely on the input image. The auxiliary prompt, automatically generated by VLM during inference, includes semantic captions and defect descriptions, indicated by <blue> and <yellow> tags, respectively.
@@ 5.2 Quantitative and Qualitative Results
Table 1 illustrates the comparative analysis of image restoration and editing techniques, evaluated via LPIPS and ManIQA metrics. The expert model – Diff-Plugin shows limited but notable performance in low-light enhancement (LPIPS/ManIQA: 0.227/0.453) and desnowing (0.133/0.508). Among generalist methods, AirNet demonstrates balanced capabilities in tasks like desnowing and dehazing, achieving LPIPS/ManIQA scores of 0.245/0.589 and 0.039/0.780, respectively. However, the instruction-driven diffusion methods reveal a more nuanced picture, with PromptFix emerging as particularly promising. It excels in colorization (LPIPS/ManIQA: 0.233/0.489), object removal (0.054/0.810), and watermark removal (0.071/0.811), consistently outperforming others. InstructP2P and InstructDiff also perform well in specific tasks, such as low-light enhancement and dehazing, but do not match the overall versatility of PromptFix. MGIE, though effective in certain domains, lacks the consistency seen in "PromptFix (Ours)," which highlights its robustness and superior performance across diverse image processing tasks. This indicates the potential of PromptFix to set new benchmarks in the field, driven by advanced instruction-driven diffusion methodologies.
Table 2: Quantitative comparison of multi-task processing on $200$ sampled test images, each paired with a degraded version exhibiting three defects (e.g., desnow, dehazy and super resolution) and the corresponding ground truth.

### v2

@@ Abstract
Diffusion models equipped with language models demonstrate excellent controllability in image generation tasks, allowing image processing to adhere to human instructions. However, the lack of diverse instruction-following data hampers the development of models that effectively recognize and execute user-customized instructions, particularly in low-level tasks. Moreover, the stochastic nature of the diffusion process leads to deficiencies in image generation or editing tasks that require the detailed preservation of the generated images. To address these limitations, we propose PromptFix, a comprehensive framework that enables diffusion models to follow human instructions to perform a wide variety of image-processing tasks. First, we construct a large-scale instruction-following dataset that covers comprehensive image-processing tasks, including low-level tasks, image editing, and object creation. Next, we propose a high-frequency guidance sampling method to explicitly control the denoising process and preserve high-frequency details in unprocessed areas. Finally, we design an auxiliary prompting adapter, utilizing Vision-Language Models (VLMs) to enhance text prompts and improve the model’s task generalization. Experimental results show that PromptFix outperforms previous methods in various image-processing tasks. Our proposed model also achieves comparable inference efficiency with these baseline models and exhibits superior zero-shot capabilities in blind restoration and combination tasks. The dataset and code are available at https://www.yongshengyu.com/PromptFix-Page.
Figure 1: We propose PromptFix, a unified diffusion model capable of performing multiple image-processing tasks. It can understand user-customized editing instructions and perform the corresponding tasks with high quality. One of the key advantages of PromptFix is high-frequency information preservation, ensuring that image details are maintained throughout VAE decoding. PromptFix can handle various images with different aspect ratios.
@@ 1 Introduction
In recent years, diffusion models [17, 55, 64] have achieved remarkable advancements in text-to-image generation. Benefiting from large-scale training on image-text pairs [57], these models can generate highly realistic and diverse images that align with text prompts. They have been successfully applied to various real-world applications, including visual design, photography, digital art, and the film industry. In addition, models trained with instruction-following data [7] have shown promising results in understanding human instruction and performing the corresponding image-processing tasks. Previous studies [19, 20, 21, 72, 71] have illustrated that with instruction-following data, we can simply fine-tune a text-to-image generation model to perform various vision tasks such as image editing [7, 19, 72], object detection [20], segmentation [21], inpainting [21, 71], and depth estimation [20]. To follow the success of these methods, we train our model utilizing input-goal-instruction triplet data for low-level image-processing tasks.
@@
We first overcome the challenge of lacking the instruction-following data for low-level tasks. Specifically, we collect image pairs by generating degraded images from the source images and adopting data from existing datasets. Then, we employ GPT4 [49] to generate the diverse text instructions for each task. We obtain $\sim$ 1.01 million input-goal-instruction triplets in the collected dataset. This dataset covers various low-level tasks including image inpainting, object creation, image dehazing, colorization, super-resolution, low light enhancement, snow removal, and watermark removal. We enrich the dataset through back-translation augmentation by swapping the inpainted and original images within the triplet and inverting the semantic orientation of the prompt. This technique effortlessly converts datasets from object removal to object creation. We also provide comprehensive details in Section 3.
@@
With the dataset, we design a new diffusion-based model named PromptFix that can understand user-customized instructions and perform the corresponding low-level image-processing tasks. In PromptFix, we address several challenges that compromise the performance of the model. First, the use of stable-diffusion architecture [55] as a generative prior often faces the issue of spatial information loss, which is caused by VAE compression [18]. Unlike unconditional or text-to-image generation, maintaining spatial detail consistency in image processing poses significant challenges, particularly with high-frequency components like text, as shown in Figure 5. To tackle this problem, we introduce High-frequency Guidance Sampling, in which we use a low-pass filter operator [48, 51] to calculate the fidelity constraint and integrate VAE skip-connect features during inference with a lightweight LoRA [25] fusion. Second, since the generative prior is not trained on low-level images, so relying solely on instructions may not always yield the desired outcome, especially when the image degradation is severe. To tackle this degradation adaptation problem, we introduce an auxiliary prompt module to provide models with more descriptive text prompts to enhance controllability for image generation. The auxiliary text prompt can be obtained by VLMs  [41]. This approach introduces a semantic caption for a degraded image and the description of its defects, such as blurriness or insufficient lighting. The auxiliary prompt module is implemented by an additional attention layer in diffusion U-Net that adapts both instruction and auxiliary prompts as conditions and intermittently omits instructional prompts during training. We identify three key advantages of this approach: 1) enabling the model to process images with severe degradation, such as extremely low-resolution images, 2) adapting the model for blind restoration for different types of image degradation, and 3) providing additional pathways for a more precise semantic representation of the target image.
@@
  • We propose a comprehensive dataset tailored for seven image processing tasks. The dataset contains $\sim$ 1.01 million diverse paired input-output images along with corresponding image editing instructions.
@@ 2.1 Instruction-guided Image Editing
Instruction-guided image editing significantly improves the ease and precision of visual manipulations by adhering to human directions. In traditional image editing, models primarily focused on singular tasks such as style transfer or domain adaptation [22, 53], leveraging various techniques to encode images into a manipulatable latent space, such as those used by StyleGAN [32]. Concurrently, the advent of text-to-image diffusion models [24, 55, 61, 63] has broadened the scope of image editing [7, 23]. Kim et al. [33] showed how to perform global changes, whereas Avrahami et al. [4] successfully performed local manipulations using user-provided masks for guidance. While most works that require only text (i.e., no masks) are limited to global editing [15, 34]. Bar-Tal et al. [6] proposed a text-based localized editing technique without using any mask, showing impressive results. For local image editing, precise manipulations are possible by inpainting designated areas using either user-provided or algorithmically predicted masks [14], all while preserving the visual integrity of the adjacent areas. In contrast, instruction-based image editing operates through direct commands like “add fireworks to the sky,” avoiding the need for detailed descriptions or regional masks. Recent approaches utilize synthetic input-goal-instruction triples [7] and incorporate human feedback [74] to execute editing instructions effectively. Despite the advances in using diffusion models for various instruction-guided image editing tasks, there is still a notable gap in research specifically addressing instruction-guided image restoration with these models. Our study aims to bridge this gap by collecting a comprehensive dataset of paired low-level instruction-driven image editing examples and proposing an all-in-one model for low-level tasks and editing.
@@ 2.2 Large Language Models for Vision
Recent advancements in the development of Large Language Models (LLMs) have led to the emergence of powerful models with extensive capabilities [13, 29, 37, 41, 75]. These LLMs, pre-trained on large-scale internet-based datasets, are equipped with broad knowledge bases that enhance their zero-shot and in-context learning abilities [5, 27, 49]. Furthermore, there is a growing focus on using LLMs for multimodal tasks [3, 28, 40, 43], incorporating methods like vision-language alignment and adapter fine-tuning. These techniques ensure that the visual data processed by visual encoders is semantically aligned with the textual input of LLMs [41]. This approach has spurred significant advancements in Text-to-Image generation, leading to the development of various LLM-based diffusion models for these tasks [26, 38, 39]. Despite these successes, there remains a relative scarcity of research focusing on using large Vision-Language Models (VLMs) for instructed image editing, particularly in detailed, low-level editing tasks.
@@ 3 Data Curation
Current off-the-shelf image datasets [7, 72, 74] with instructional annotations primarily facilitate image editing research, encompassing tasks such as color transfer, object replacement, object removal, background alteration, and style transfer. Nevertheless, their overlap with low-level applications is limited. Moreover, we find it challenging to achieve satisfactory results for existing models in image restoration. Our goal is to construct a comprehensive visual instruction-following dataset specifically for low-level tasks. We obtain $\sim$ 1.01 million training triplet instances.
Paired Image Collection. We initially gather source images from various existing datasets. Subsequently, we produce degraded and inpainted images to create an extensive set of paired image data. We compile approximately two million raw data points across eight tasks: image inpainting, object creation, image dehazing, image colorization, super-resolution, low-light enhancement, snow removal, and watermark removal. For the test set, we randomly select 300 image pairs for each task. More details about the dataset composition are provided in the Appendix A.1.
@@ 4.1 Diffusion Model
where $\alpha_{t}$ and $\sigma_{t}$ are coefficients that manage the signal-to-noise ratio at each interpolation point $\mathbf{z}_{t}$. This process aims to maintain variance, adopting coefficient strategies as detailed in sources such as [31]. Modeled as a stochastic differential equation (SDE) in continuous time, the forward process can be expressed as $d\mathbf{z}=\mathbf{f}(\mathbf{z},t)dt+g(t)d\mathbf{w_{t}},$ where $\mathbf{f}(\mathbf{z},t)$ is a vector-valued drift coefficient, $g(t)$ is the diffusion coefficient, and $\mathbf{w_{t}}$ represents Brownian motion at time $t$.
@@
During the training phase, a simple diffusion loss [24] is utilized, whereby the neural network still employs forward inference to predict noise. The sample data estimate $\mathbf{z}_{0}$ can be obtained at any step $t$ by using the current noisy data and the predicted noise and is derived as:
@@ 4.3 High-frequency Guidance Sampling
There is a fundamental requirement in image restoration and generation tasks: the processed image must maintain high accuracy in semantics. We observe that vanilla VAE reconstructions tend to lose image details such as textual rendering, which contains high-frequency information, as shown in Figure 5. Therefore, we propose high-frequency guidance sampling to balance the quality and fidelity of generation.
@@
Table 1: Quantitative comparison is conducted across seven low-level datasets with a $512\times 512$ input resolution. Expert models refer to approaches, such as Diff-plugin [44], which use non-generalizable training pipelines and maintain separate pre-trained weights for each of the four restoration tasks. Image Restoration Generalist Methods denote models that integrate multiple low-level tasks into a single framework. Instruct-driven Diffusion Methods represent diffusion generative baselines that follow human language instructions. $\uparrow$ indicates higher is better and $\downarrow$ indicates lower is better. The best and second best results are in bold and underlined, respectively.
@@
We propose a fidelity constraint to model the spatial discrepancy between the image and the ground truth. We implement two types of high-pass operators to extract high-frequency signals from the degraded image. For the Fourier filtering operator $\mathcal{F}(\cdot)$, we convert the generated image from the spatial to the frequency domain using the Discrete Fourier Transform [48]. High-frequency components are then isolated via high-pass filtering and reconstituted into an image through the Inverse Fourier Transform [48]. Meanwhile, we apply the Sobel [51] edge detection operator $\mathcal{S}(\cdot)$ as a complement. The fidelity constraint evaluates the divergence between the high-frequency components of the ground truth and the processed image, ensuring the preservation of spatial information throughout the sampling process. Additionally, to obtain the image at time step $t$, we utilize predicted noise $\epsilon$ from diffusion model $\mathcal{H}$ to compute the $\mathbf{z}_{t\rightarrow 0}$ estimation at any time step. The fidelity constraint is calculated as follows:
@@
Given that $\mathbf{z}_{t>0}$ represents a noisy latent variable, assigning equal weight to each timestep’s latent is impractical. To mitigate the cumulative error induced by this practice, we introduce a time-scale weight $e^{-\lambda t}$. The overall sampling algorithm is detailed in Algorithm 4.1.
@@ 5.1 Experimental Setup
Baselines and metrics. We adopt instruction-based generalist models, such as InstructP2P [7], MGIE [19], and InstructDiffusion [21], as our primary comparison. MGIE employs VLM-guided techniques for image editing, while InstructDiffusion addresses overlapping tasks with our training objectives, including watermark removal and inpainting. Additionally, we evaluate all-in-one image restoration methods like AirNet [36] and PromptIR [50] (which do not support instruction input), as well as image restoration expert models, fine-tuned for specific sub-tasks [44, 71]. We assess the similarity of the generated images to the ground truth using metrics such as PSNR, SSIM [67], and LPIPS [73]. For no-reference image quality evaluation, we utilize the ManIQA [70] metric.
Figure 4: Qualitative analysis of VLM-guided blind restoration for desnowing, dehazing, and low-light enhancement. The results are obtained from PromptFix without explicit task instructions, relying solely on the input image. The auxiliary prompt, automatically generated by a VLM during inference, includes semantic captions and defect descriptions, indicated by <blue> and <yellow> tags, respectively.
@@ 5.2 Quantitative and Qualitative Results
Table 1 illustrates the comparative analysis of image restoration and editing techniques, evaluated via LPIPS and ManIQA metrics. The expert model – Diff-Plugin shows limited but notable performance in low-light enhancement (LPIPS/ManIQA: 0.227/0.453) and desnowing (0.133/0.508). Among generalist methods, AirNet demonstrates balanced capabilities in tasks like desnowing and dehazing, achieving LPIPS/ManIQA scores of 0.245/0.589 and 0.039/0.780, respectively. However, the instruction-driven diffusion methods reveal a more nuanced picture, with PromptFix emerging as particularly promising. It excels in colorization (LPIPS/ManIQA: 0.233/0.489), object removal (0.054/0.810), and watermark removal (0.071/0.811), consistently outperforming others. InstructP2P and InstructDiff also perform well in specific tasks, such as low-light enhancement and dehazing, but do not match the overall versatility of PromptFix. MGIE, though effective in certain domains, lacks the consistency seen in "PromptFix (Ours)." This highlights the robustness and superior performance of PromptFix across diverse image processing tasks and indicates the potential of PromptFix to set new benchmarks in the field, driven by advanced instruction-driven diffusion methodologies.
Table 2: Quantitative comparison of multi-task processing on $200$ sampled test images, each paired with a degraded version exhibiting three defects (e.g., desnowing, dehazing and super-resolution) and the corresponding ground truth.
@@
Figure 3 illustrates the visual comparison across all selected baseline models. In colorization, our PromptFix produces the most visually accurate and vibrant results, closely resembling the ground truth. For the watermark removal task, it effectively restores the original image without introducing artifacts, outperforming MGIE [19] and other methods. In desnowing and low-light enhancement, PromptFix achieves clearer and more natural outputs, significantly reducing noise and enhancing visibility. Additionally, in super-resolution, PromptFix demonstrates remarkable clarity and accuracy, preserving fine details and surpassing all comparative methods. In dehazing, although PromptFix’s performance is visually comparable to image restoration experts PromptIR [50] and AirNet [36], PromptFix outperforms the recent stable-diffusion-based method Diff-plugin [44], achieving a clean, sharp appearance, and closely matching the ground truth.
5.3 Ablation Study
Effectiveness of High-frequency Guidance Sampling. The High-frequency Guidance Sampling (HGS) method is introduced to balance fidelity and quality. To validate the effectiveness of HGS, we conduct qualitative experiments and quantitative experiments. As depicted in Figure 5, in a low-light scenario, the model aims to enhance the visibility (quality) of the input image while preserving its original textual details (fidelity). For baselines leveraging stable-diffusion as a generative prior, the strong compression capability of VAE also brings the issue of spatial information loss, as demonstrated by InstructDiff [21], MGIE [19], and Diff-Plugin [44] in Figure 5. This issue is independent of the model’s ability to effectively follow instructions. As shown by the variant “Ours w/o HGS,” our method significantly enhances low-light images compared to the three baselines, yet still fails to retain small-scale textual structures. By incorporating HGS, as seen in “Ours,” the proposed framework delivers a high-fidelity solution that also meets the low-light enhancement instruction. The usage of $\mathcal{F}(\cdot)$ and $\mathcal{S}(\cdot)$ improves the quality of the image generated, demonstrated by the quantitative results shown in Table 5.